OpenStudy (bambimonster):

1.in true false examination with 5 questions, the probability that an unprepared student will score all correct by chance is ? 2. a manufacturer produces a certain kind of cereal. suppose the weight of the box of their product is normally distributed with mean 500g and std.dev 10g. a customer bought 3 boxes of this type of cereal. the probability that among 3 boxes, two of them contain less than 500g is? 3.the binomial distribution is symmetric when p is equal to? a. 0.25 b.0.1 c.0.5 d.not(1-p) e. any value 4. the probability of observing no fewer than two heads in four tosses of a fair coin is?

OpenStudy (anonymous):

#1. 0.5*0.5*0.5*0.5*0.5=0.5^5=0.03125 or 3.125%

OpenStudy (bambimonster):

hey how did u get 0.5?

OpenStudy (anonymous):

2. The std dev is irrelevant. P(2 are less than 500g) = C(3,2) x 0.5^3 = 3 x 0.125 = 0.375

OpenStudy (bambimonster):

a new company receives a very large shipment of computer chips from a discount broker. unknown to the company, 10% of the computer chips are defective. the company decides to accept the shipment if fewer than two of 20 randomly selected computer chips are found to be defective. the probability that the shipment is accepted is?

OpenStudy (bambimonster):

C(3,2) x 0.5^3 i dont understand this part

OpenStudy (cwrw238):

its 0.5 or 1/2 because the probability of choosing 1 from the 2 possibilities in each question ( true or false) is one out of two

OpenStudy (bambimonster):

ok thanks!

OpenStudy (bambimonster):

can u help me with no.4?

OpenStudy (anonymous):

this one? the probability of observing no fewer than two heads in four tosses of a fair coin is?

OpenStudy (anonymous):

"no fewer than two heads" means 2, 3, or 4 " there are \(2^4=16\) possible outcomes one way to do this is by counting

OpenStudy (anonymous):

or just note that there is one way to get no head (h,h,h,h) and 4 ways to get only one head (h, t, t, t) (t, h , t, t) (t, t, h , t) (t, t ,t, h) there fore that are \(16-5=11\) ways to get two or more heads

OpenStudy (anonymous):

lol one way to get no heads should have been (t, t, t, t)

OpenStudy (bambimonster):

hey thanks but the answer is 0.6875

OpenStudy (bambimonster):

@satellite73